OpenStudy (anonymous):

GUYs... quick two ice cream trucks will visit Jeannette's neighborhood in 20 days and in 40 days. However, the problem asks: when is the next time both trucks will visit on the same day?

OpenStudy (anonymous):

I get the answer for you, If I deserve medal can you give me one?

OpenStudy (anonymous):

Take me so long to think

OpenStudy (anonymous):

sry was not on this page... answer is 20 :) :)

OpenStudy (anonymous):

ok Nancy

OpenStudy (anonymous):

what this you say?

OpenStudy (anonymous):

Did you know the answer?

OpenStudy (anonymous):

you say the key is 20 time?

OpenStudy (anonymous):

here the answer: let x for 20 days , y for 40 days x+y=60 x=60-y 20x=40y Subtitution x=60-y 20(60-y)=40y 120 - 20 y = 40y 120 = 40y+20y 120 = 60y 120/60=y 2 = y Sutitution y= 2 solve for y 20x=40(2) 20x =80 x = 80/2 x = 4 Check: 20(4) = 40(2) 80 = 80 The truck 20 days trucks will visit 4 time The truck 40 days trucks will visit 2 time If trucks 20 days trucks will visit 20 time than trucks 40 days trucks will visit 40 time or you can twice time the truck 20 day will equal to the trucks 40 days

OpenStudy (anonymous):

good cluck your home work